OpenStudy (anonymous):

Will fan and give medal to best answer!!!!! 1.5 x 106 is a number written in scientific notation. In order to convert this number to standard notation, we need to: A. Divide the number by six. B. Move the decimal point six places to the right. C. Move the decimal point six places to the left. D. Multiply the number by six.

OpenStudy (anonymous):

Scientific notation works by dividing a number by 10 until it only has a ones digit (along with some decimal values) left over. The number of times it was divided by 10 is then used as the exponent over 10. For example:\[4200 = 420 * 10^1 = 42 * 10^2 = 4.2 * 10^3\]So, do you think the original number was very large or very small?